# Numerical Homogenization by Continuous Super-Resolution

### v1

@@ IV-C Ablation studies and analysis
Gabor based grid sampling. We are also interested in the effect of using Gabor-based grid sampling for the coordinate encoding. To validate its effect, we compare it with different encoding schemes, including plain encoding without learnable parameters, sinusoid [38], IPE [35], and LTE [56]. In Table II, we can see that our method achieves the best performance in terms of MSE, SSIM, and MAE.
@@
Figure 8: Visualization of multi-scale feature sampling for 16-times super-resolution. Given one 16-times super-resolved example ((a) shows the ground truth data), we visualize the frequency map of coarse- and fine-scale reconstructed data in (b), and the plots of sliced pixels in (c). The slicing position is shown with the red lines.
TABLE III: Ablation study on the multi-scale feature sampling and SCS loss. We report the MSE ($10^{-3}$), MAE ($10^{-3}$) and SSIM results on the test datasets of 4-times and 8-times upsampling.
Multi-scale 4x 8x
@@
Single-scale+$\mathrm{L}_{1}$ 0.867 0.858 21.46 0.946 0.922 18.12
Single-scale+$\mathrm{L}_{1}$+$\mathrm{L}_{\mathrm{scs}}$ 0.857 0.860 21.15 0.942 0.923 17.86
Multi-scale+$\mathrm{L}_{1}$ 0.848 0.859 20.86 0.907 0.920 17.45
Multi-scale+$\mathrm{L}_{1}$+$\mathrm{L}_{\mathrm{scs}}$ 0.834 0.863 20.09 0.885 0.924 16.93
Multi-scale feature sampling and loss optimization. We conduct an ablation study to show the effect of multi-scale feature sampling and the SCS loss for super-resolution. In Table III, we mark single-scale as the model that only uses coarse-level features for reconstruction (first equation in (9)), and multi-scale as the model that uses both coarse- and fine-level features for reconstruction (first and second equations in (9)). LIIF is the baseline for comparison and we can see that using both multi-scale feature sampling and the novel SCS loss can improve the reconstruction quality in terms of MSE and SSIM. To further show the effect of using SCS loss, we also plot the validation loss curve in Figure 9. We can see that using the SCS loss, the validation loss is lower and it also converges faster than without using the SCS loss.
@@
To demonstrate the efficiency of multi-scale feature sampling, we show one example of 16-times super-resolution in Figure 8. In (b), we apply the fast Fourier transform (FFT) to the super-resolved data and keep the frequency energy spectrum, and then we shift the zero-frequency component to the center of the array. We can observe that the coarse-scale reconstructed image focuses more on the low-frequency estimation, while the fine-scale reconstruction expands the energy to the wider range of other frequency bands, which indicates that it learns to recover higher-frequency information. In Figure 8 (c), we plot the pixel values in the sliced regions (marked by the red lines in (a)). We use the solid dark brown line to represent the results obtained by our final model (combined coarse- and fine-scale feature sampling), and the solid yellow line to represent the results obtained only by coarse-scale feature sampling. We can see that the final result is closer to the ground truth which aligns with the observation in (b) that more details are recovered by the multi-scale feature reconstruction. The alignment is still not fully optimal but recall that we are dealing with a complicated setting far beyond the training data. Including further coefficient maps into the training process, can further improve the results.
@@
Conditional super-resolution. The proposed NH-CSR approach takes a coefficient map as the conditional input to jointly learn the super-resolved data. To demonstrate its effect, we provide two examples in Figure 10. Comparing the first and the third column, we may conclude that using coefficient maps as conditions for super-resolution can significantly improve the visual similarity. From the second, third, and fourth row, it becomes clear that using a more complex global feature encoder can improve the reconstruction quality. It is worth noting that using swin transformers [50] (cf. the fourth column) as the encoder would cost much more computation time than RDN [65], and it also produces higher MSE losses.

### v2

@@ IV-C Ablation studies and analysis
Gabor based grid sampling. We are also interested in the effect of using Gabor-based grid sampling for the coordinate encoding. To validate its effect, we compare it with different encoding schemes, including plain encoding without learnable parameters, sinusoid [siren], IPE [NERF], and LTE [lte]. In Table II, we can see that our method achieves the best performance in terms of MSE, SSIM, and MAE.
@@
Figure 8: Visualization of multiscale feature sampling for 16-times super-resolution. Given one 16-times super-resolved example ((a) shows the ground truth data), we visualize the frequency map of coarse- and fine-scale reconstructed data in (b), and the plots of sliced pixels in (c). The slicing position is shown with the red lines.
TABLE III: Ablation study on the multiscale feature sampling and SCS loss. We report the MSE ($10^{-3}$), MAE ($10^{-3}$) and SSIM results on the test datasets of 4-times and 8-times upsampling.
Multiscale 4x 8x
@@
[…]
Multiscale feature sampling and loss optimization. We conduct an ablation study to show the effect of multiscale feature sampling and the SCS loss for super-resolution. In Table III, we mark single-scale as the model that only uses coarse-level features for reconstruction (first equation in (9)), and multiscale as the model that uses both coarse- and fine-level features for reconstruction (first and second equations in (9)). LIIF is the baseline for comparison and we can see that using both multiscale feature sampling and the novel SCS loss can improve the reconstruction quality in terms of MSE and SSIM. To further show the effect of using SCS loss, we also plot the validation loss curve in Figure 9. We can see that using the SCS loss, the validation loss is lower and it also converges faster than without using the SCS loss.
@@
To demonstrate the efficiency of multiscale feature sampling, we show one example of 16-times super-resolution in Figure 8. In (b), we apply the fast Fourier transform (FFT) to the super-resolved data and keep the frequency energy spectrum, and then we shift the zero-frequency component to the center of the array. We can observe that the coarse-scale reconstructed image focuses more on the low-frequency estimation, while the fine-scale reconstruction expands the energy to the wider range of other frequency bands, which indicates that it learns to recover higher-frequency information. In Figure 8 (c), we plot the pixel values in the sliced regions (marked by the red lines in (a)). We use the solid dark brown line to represent the results obtained by our final model (combined coarse- and fine-scale feature sampling), and the solid yellow line to represent the results obtained only by coarse-scale feature sampling. We can see that the final result is closer to the ground truth which aligns with the observation in (b) that more details are recovered by the multiscale feature reconstruction. The alignment is still not fully optimal but recall that we are dealing with a complicated setting far beyond the training data. Including further coefficient maps into the training process, can further improve the results.
@@
Conditional super-resolution. The proposed NH-CSR approach takes a coefficient map as the conditional input to jointly learn the super-resolved data. To demonstrate its effect, we provide two examples in Figure 10. Comparing the first and the third column, we may conclude that using coefficient maps as conditions for super-resolution can significantly improve the visual similarity. From the second, third, and fourth row, it becomes clear that using a more complex global feature encoder can improve the reconstruction quality. It is worth noting that using swin transformers [swinir] (cf. the fourth column) as the encoder would cost much more computation time than RDN [rdn], and it also produces higher MSE losses.
Figure 11: Visualization of real-world soil patterns for 16-times super-resolution. To demonstrate the effect of the proposed method for real-world data, we visualize the super-resolved finite element data. For better comparison, the values are multiplied by a factor of 10 and taken modulo 255 as before.
IV-D Real-world applications.
We are interested in applying our proposed method to real-world data to validate its robustness without domain-specific model fine-tuning. In Figure 11, we show comparisons on two datasets inspired by two research works: 1) Ye et al. (2024) [YeXZC24] and 2) Vaz et al. (2014) [VazTLC14]. The first row shows the binary microCT soil patterns used as inputs to our FEM solver to generate paired low-resolution fields ($33\times 33$) and high-resolution references (fourth row, $513\times 513$). All competing methods are tasked with performing $16$-times super-resolution. Across both datasets, the bicubic baseline produces overly smoothed predictions with little structural detail, while LIT partially restores coarse patterns but still lacks sharp transitions (introducing extra ringing effects) and accurate spatial variability. In contrast, our method (fourth row) recovers markedly clearer boundaries, more faithful spatial gradients, and finer heterogeneous structures that are visually much closer to the ground truth. These improvements are especially evident in regions with sharp contrast or irregular soil clusters, where competing methods tend to oversmooth or distort the shapes. This demonstrates that our approach generalizes well to real-world data and preserves physically meaningful structures even without task-specific retraining.
@@ Acknowledgments
VI Biography Section
[Uncaptioned image] Zhi-Song Liu received his Ph.D. (2016-2020) from The Hong Kong Polytechnic University, Artificial Intelligence and Signal Processing Lab. Previously, he was a principal research scientist at DELL EMC. Before that, he was a research fellow (01/2021-12/2021) in CIHE and a post-doc researcher in LIX, Ecole Polytechnique (01/2020-12/2020). His research interests include image and video signal processing, computer vision, deep learning, and 3D data processing. He joined the CVPR Lab at LUT University in 2023 as an associate professor.
[Uncaptioned image] Roland Maier studied Mathematics at the University of Bonn (2012-2017) and received his Ph.D. from the University of Augsburg in 2020. After some time as a postdoc at the University of Gothenburg and Chalmers University of Technology, he became a junior professor at University of Jena in 2021. In 2023, he joined Karlsruhe Institute of Technology as a tenure-track professor for numerics of partial differential equations. He mainly works on multi-scale methods and numerical homogenization approaches.
[Uncaptioned image] Andreas Rupp received his M.Sc. and PhD degrees from the FAU Erlangen–Nuremberg in December 2015 and May 2019, respectively. He was a postdoctoral researcher at Heidelberg University till June 2021. Since then, he was a tenure-track assistant professor at LUT University and, since September 2023, an Academy of Finland research fellow. In September 2024, he joined Saarland University as a professor of applied mathematics.